OpenStudy (anonymous):

how can i work out tan(19pi/4) without a calculator...???

OpenStudy (anonymous):

\[\frac{19}{4}=4\tfrac{3}{4}\] and so \[\frac{19\pi}{4}=2\pi+\frac{3\pi}{4}\]

OpenStudy (anonymous):

since tangent is periodic with period pi, you can find this number by finding \[\tan(\frac{3\pi}{4})\] which if you don't know it amounts to finding \[\sin(\frac{3\pi}{4})\] \[\cos(\frac{3\pi}{4})\] and then dividing the first by the second

OpenStudy (anonymous):

kinda get that... but then... how do you find sin(3pi/4)....??